Zeronknight (zeronknight):

Choose the value that is a solution to the following inequality. 7x - 9 \leq 40 A. 7 B. 9 C. 10 D. 8

Zeronknight (zeronknight):

\[7x - 9 \leq 40\]

OpenStudy (dayakar):

add both sides 9 simplify

Zeronknight (zeronknight):

add 9 to both sides? then divide 7 and 40?

Zeronknight (zeronknight):

you can't simplify 7 by 2

OpenStudy (dayakar):

\[7x-9+9\le40+9\]

OpenStudy (dayakar):

\[7x \le49\] now divide both sides with 7

OpenStudy (dayakar):

can u do it

Zeronknight (zeronknight):

yeah

Zeronknight (zeronknight):

7 / 7 = 1 49 / 7 = 7

OpenStudy (dayakar):

\[x \le7\]

OpenStudy (dayakar):

therefore x=7 or x<7

OpenStudy (dayakar):

do u get it

Zeronknight (zeronknight):

X < 7 Cause X is equal to 1 basically

OpenStudy (dayakar):

it is an inequation it has many solutions if u choose x=7 it is true and if u choose any value less than 7 because x <7 it is true

OpenStudy (dayakar):

in our problem we take x=7

Zeronknight (zeronknight):

oh

Zeronknight (zeronknight):

so it's A?

OpenStudy (dayakar):

yes

OpenStudy (dayakar):

yes

Zeronknight (zeronknight):

Ok thank you!